Saturday 03/03/12

11:00A – 7:00 P

  • Overnight run of Psc H3K27 aborted after ~23 sections (235 Gb) .
  • Moving files from Alistair2 (G:\) to Alistair3 (K:\) to mobilize Alistair2 for data collection tonight.
  • Can copy files over local intranet quickly from G drive but not D drive on Zidane.
  • Getting chromatic correction via bead slides to work.
  • check with Josh and Yari about Git code sharing.
  • Initialized local git repo for General_STORM\Code
  • Analysis of high threshold H2 data.
  • Goal: share an external hard-drive from Zidane, have Tuck connect to hard-drive through ethernet, poll hard-drive for existence of data files (as they are written) and start analysis as data appears.
  • 1 hour with shaking seems to dissolved a bit of the sections for Hb hb(y) and Psc K27me3.
  • working on Fly Meeting presentation.
  • working on chromatic alignment of S2 images.
  • STORM2 alignment off. ┬áCan not image.
This entry was posted in Summaries and tagged , . Bookmark the permalink.